OpenStudy (anonymous):

SOLVE THE UNKNOWN N/4-7=8

zepdrix (zepdrix):

Hey angie! \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\) Is this what the problem looks like?\[\Large \frac{N}{4}-7\quad=\quad 8\]

OpenStudy (anonymous):

YEA

zepdrix (zepdrix):

Hmm so we want to try to get our N variable alone by itself. Let's start by `adding` 7 to each side.\[\Large \frac{N}{4}-7\color{red}{+7}\quad=\quad 8\color{red}{+7}\]What does that do to the 7's on the left side of the equation?

OpenStudy (anonymous):

NOT SURE

zepdrix (zepdrix):

So on the left side we're subtracting 7 and adding 7. That part a little confusing? hmm

zepdrix (zepdrix):

Think about this example really quick: So let's say we have 3 cookies. If I subtract 1 cookie, we have 2 cookies. If I then add 1 cookie, we have 3 cookies again right? What we did was, 3-1+1=3 The minus 1 and plus 1 give us nothing when we combine them.

OpenStudy (anonymous):

THIS IS MY FIRST TIME DOING THIS

zepdrix (zepdrix):

ok :)

zepdrix (zepdrix):

So let's just try to get through it, and see which parts are most confusing.\[\Large \frac{N}{4}-7+7\quad=\quad8+7\]So the -7 and +7 can combine to give us zero.\[\Large \frac{N}{4}\quad=\quad 8+7\]

zepdrix (zepdrix):

What does 8 + 7 give you? :)

OpenStudy (anonymous):

15

zepdrix (zepdrix):

\[\Large \frac{N}{4}\quad=\quad15\]Ok good.

OpenStudy (anonymous):

THANK YOU

zepdrix (zepdrix):

To further solve for N, we need to first recognize that it's being `divided` by 4. So to move the 4 to the other side, we must perform the inverse operation of `division`, which is `multiplication`.

zepdrix (zepdrix):

Multiplying both sides by 4 gives us,\[\Large 4\cdot \frac{N}{4}\quad=\quad 15\cdot 4\]We have a 4 being multiplied `and` divided on the left side, so they can "cancel out".\[\Large \cancel4\cdot\frac{N}{\cancel4}\quad=\quad15\cdot4\]Which simplifies to,\[\Large N\quad=\quad 15\cdot4\]

zepdrix (zepdrix):

What does 15 times 4 give you? :o

OpenStudy (anonymous):

60

zepdrix (zepdrix):

Good. So that would be our final answer.\[\Large N\quad=\quad 60\]

zepdrix (zepdrix):

I know it's a little tricky to get through :( Ask questions if you're confused about any specific parts.

OpenStudy (anonymous):

THANK YOU SO MUCH